mykenzie:

Circle A has a diameter of 9 inches, a circumference of 28.26 inches, and an area of 63.585 square inches. The diameter of circle B is 5 inches, the circumference is 15.70 inches, and the area is 19.625 square inches.. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

AZ:

Part A) Circumference = 2*\(\pi\)*r = diameter * \(\pi\) and they told you circle A has a diameter of 9 inches and circumference of 28.26 so C = D* pi plug in C = 28.26 and D = 9 and find the value of \(\pi\) and do the same thing for circle B which has a diameter of 5 inches and a circumference of 15.70 and compare the two values of \(\pi\) that you get